Status Report

NASA Space Station On-Orbit Status 10 Apr 2004

By SpaceRef Editor
April 10, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 10 Apr 2004

All ISS systems continue to function nominally, except those noted previously or below.   Saturday — light-duty day for the crew, and their 173rd day onboard ISS (175 days in space).  Happy Easter to all!

As most every Saturday, the crew started the day with the weekly 3-hr. station cleaning.   [This included removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

FE Alexander Kaleri performed the daily routine maintenance of the SOZh environmental control & life support systems in the Service Module.

CDR/SO Mike Foale and Alex Kaleri worked out on their regular daily physical exercise program of 2.5 hrs on TVIS treadmill, CEVIS bike, RED exerciser and VELO cycle with load trainer.

At 9:00am EDT, the crew conducted the weekly planning conference with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

Also, at 10:45am the crew also held their monthly teleconference with ISS Program Management via S-band/audio.

Both crewmembers had their weekly PFCs (private family conferences), Sasha via VHF/home phone, Mike via Ku- plus S-band.

Working off the Russian task list, Kaleri prepared a PC hard disk (HDD-5) with all collected Uragan experiment data, copied from the camera digital memory cards.  The HDD will be returned on Soyuz 7S, and a new disk (HDD-6) will be delivered on Soyuz 8S for Uragan imagery.

At 3:00pm, Mike Foale downlinked special greetings from the ISS to be played on Cosmonautics Day, next Monday (4/12), in celebration of Yuri Gagarin’s pioneering flight 43 years ago.  This day also holds special significance as the 23rd

anniversary of the first Space Shuttle flight (SST-1) on 4/12/81 with John Young and Bob Crippen.   [“Today, Sasha and I reflect on these milestones.  Space flight is still in its infancy, but it has brought us immense benefits that enhance our lives on Earth and increase our knowledge of the universe.  The cooperation in space today between all of the partners involved in the ISS is remarkable.  That cooperation has built a program that is stronger than could have been achieved in any other fashion.”]

Yesterday, Mike Foale repaired one of three failed GLA (General Luminaire Assemblies) lighting units in the Node by replacing its LHA (light housing assembly) with an LHA cannibalized from the U.S. Airlock’s crewlock compartment.  A second GLA, which has both its LHA and BBA (baseplate ballast assembly) failed, is desired for 7S/8S joint crew ops in the Node, and its replacement with a unit scavenged from another location in the Node (behind the starboard hatch) was added to Foale’s “job jar” task list for this weekend.  The third GLA remains failed.

Update on TEPC:  Mike Foale’s inspection yesterday of the U.S. TEPC (tissue equivalent proportional counter) cables for damaged pins and sockets did not locate the cause of frequent spontaneous TEPC reboots (similar to a PC reset) ever since its deployment after Progress 13P arrival.  Afterwards, TEPC files were downloaded to the medical equipment computer (MEC) to help confirm that the file downlinks over S-band are not corrupted.

Update on PCU-2:  As of yesterday, after 13 open/close cycles to seat it properly, latch valve 1 of PCU-2 (plasma contactor unit 2) continues to leak.  Today, the backup valve, latch valve 2, was to be commanded closed.   [The impact of leaking xenon through valve 1, which is located upstream at the xenon tank, is negligible.  There is approximately 438 EVA days worth of the gas remaining in the tanks.  This particular valve has leaked twice before since PCU deployment and was successfully seated each time.  The two PCUs, 351-pound electron “guns” on the Z1 truss, “shoot” accumulated electrons away from the ISS structure, with the discharge essentially creating an electrical short to the ambient plasma.  Their emission is facilitated by the use of the inert (“noble”) gas xenon, which is ionized and expelled thermionically (i.e., by heating) via a hollow cathode assembly.  PCU operation requires considerable power of 420 W each (120 V dc @ 3.5 Amp max. current).  The units are intended to reduce the station’s “floating” potential to less than 40 V but do not themselves measure the potential or the plasma environment.]

Weekly Science Update (Expedition Eight — 21st):

GASMAP:   Mike was thanked for his work with the GASMAP this Increment.

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound (ADUM):   No more activities scheduled for this Increment.

Hand Posture Analyzer (HPA:  The ground is looking forward to the final session next week and analyzing the consolidated data set.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Sample PFMI-5A was successfully completed this week.  Two different growth rates were used.  The planar interface became dendritic after one hour and 45 minutes at a growth rate of 1 micron/sec and after 4 hours at a growth rate of 0.7 microns/sec. These data will be compared to theoretical models. Thanks to Mike for his efforts “enabling successful PFMI science during this increment”.
Space Acceleration Measurement System (SAMS):   SAMS is operating nominally. SAMS successfully supported the PFMI run on 4/5.  On 4/6, SAMS successfully powered and booted RTS/D1.  It ran with no issues on 4/6, and RTS/D2 was powered and booted nominally 4/7. SAMS will continue to closely monitor the SAMS unit on ISS.
Microgravity Acceleration Measurement System (MAMS):   MAMS-OSS continues to measure low-frequency accelerations to characterize the ISS environment.  MAMS-HiRAP is currently active and acting as a backup to SAMS.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Mike was thanked for keeping the filters clean so that temperatures could remain nominal.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):   Experiment has ended.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):   BCAT-3 thanked Mike for “an interesting and highly successful set of experiment sessions”.  The data has already confirmed the value of this microgravity research, giving researchers much useful data, and they will spend some time working to accurately interpret it.  Mike’s work has “left nothing to be desired”.

Renal Stone (RS):   The last in-flight collection and diet logging will begin next Friday (4/16).  The crew was thanked again for their support with daily pill ingestion.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   No more activities scheduled for this Increment.

Foot/Ground Reaction Forces during Space Flight (FOOT):   To crew: “Great job on the Foot Ops this week.  We look forward to reviewing the data.”

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Mike’s patience and work with this payload during this increment is being appreciated.  His contributions to the crew conferences with the PI provided great insight towards the experiment. The method he described will be used to “create a procedure for future increments to remove bubbles from the syringe and the TCM bag”.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Students are still learning from the images captured by EarthKAM during this expedition.

Earth Science Toward Exploration Research (ESTER):  No more activities scheduled for this increment.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   Nothing new.

Miscible Fluids in Microgravity (MFMG):    Four runs are remaining; they will be performed on Increment 9.

Educational Payload Operations (EPO):   The educational demonstrations performed on this increment are being edited for release and distribution to schools around the country.

Crew Earth Observations (CEO):  The crew’s outstanding image from two weeks ago of the Betsiboka River Delta in Madagascar will be published in Earth Observatory this weekend.  The article highlights the dynamic changes in the delta’s channel and island pattern due to high runoff and erosion rates associated with tropical storms rainfall and land use over the watershed.  CEO investigatorsalso just reviewed a more recent set of images taken from ISS of that area on ground request, following Tropical Cyclone Fay.  The crew was thanked for the excellent images of the Bombay region and the fine excursion through east Africa’s lakes prior to their mapping of the Omo River delta and the most detailed views of Addis Abeba the team has ever seen.  Camera times remain good.  Thanks to the crew for their continued efforts at CEO despite the limitations imposed by attitude and window access.

Today’s optional CEO targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Cairo, Egypt (excellent viewing conditions existed for this nadir pass over the Egyptian capital.  Trying to capture the whole of the city in a single frame), Nile River Delta (nice mapping pass over the middle delta region that also included the target city of Cairo), Tigris-Euphrates, Turkey (nice spring weather over this region will enhance snowmelt and provide good stream flow.  Looking for opportunities to map the water levels in the numerous new reservoirs constructed in this region), Aral Sea (ISS had a nice first pass of the day, in fair weather, over the northern periphery of this vanishing water body.  Looking right of track for possible sun glint enhancement.  On a second pass, light was not as good, but nadir mapping views of the southwestern shoreline were possible), Bamako, Mali (ISS had a nadir pass in fair weather over the Mali capital, situated on the upper Niger River), Bahamas (400mm-lens:  ISS passed over the far southeastern end of the Bahamas where the islands are small, but the banks are large.  Using the long lens in nadir views for details of the coral reef structures), and Athens, Greece (nadir pass in late afternoon light over the Greek capital).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

ISS Orbit  (as of this morning, 8:02am EDT [= epoch]):

  • Mean altitude — 364.3 km
  • Apogee — 371.6 km
  • Perigee — 357.0 km
  • Period — 91.83 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0010832
  • Orbits per 24-hr. day — 15.68
  • Mean altitude loss last 24 hours — 90 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30769

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.